OpenStudy (anonymous):

Hellllpppp :'( Solve 3^(x – 4)= 7^(x + 9)

OpenStudy (anonymous):

hey angellaaa :). You seem nice, can you help me??

OpenStudy (angela210793):

I wish I could...i did smth..but i stuck later :(( this is wht wolframalpha says though.... http://www.wolframalpha.com/input/?i=3^%28x+%E2%80%93+4%29%3D+7^%28x+%2B+9%29

sam (.sam.):

3^(x-4)=7^(x+9) ln(3^(x-4))=ln(7^(x+9)) xln(3)-4ln(3)=ln(7^(x+9)) xln(3)-4ln(3)-xln(7)=9ln(7) x(ln(3)-ln(7))=4ln(3)+9ln(7) x=(4ln(3)+9ln(7)) / (ln(3)-ln(7))

OpenStudy (angela210793):

Thanks Sam :D

sam (.sam.):

no problem :)

OpenStudy (anonymous):

what the answer though ? heheh didnt get it lol

OpenStudy (anonymous):

did you get it agela?

OpenStudy (anonymous):

\[\ln(3^{x-4)}=\ln(7^{x+9)}\]

OpenStudy (anonymous):

xln3-4ln3=xln7+9ln7

OpenStudy (anonymous):

x(ln(3)-ln(7))=4ln(3)+9ln(7)

OpenStudy (anonymous):

x(1.1-1.9)=4.39+17.51 x(-0.8)=21.9 x=21.9/-0.8 x=-27.4

OpenStudy (anonymous):

if my opptions are: x ≈ –0.04 x ≈ –2.59 x ≈ –25.86 x ≈ –0.39 which one is it??? :S

OpenStudy (anonymous):

c!! since its close to it..

OpenStudy (anonymous):

waw thnks mattthh

OpenStudy (anonymous):

no problem, :]